# The Modularity of an Abelian Variety

### v1

@@ Abstract.
We introduce the concept of the modularity of an abelian variety defined over the rational number field extending the modularity of an elliptic curve. We discuss the modularity of an abelian variety over $\mathbb{Q}$. We conjecture that a simple abelian variety over $\mathbb{Q}$ is modular. The detailed article will appear later.
Subject Classification: Primary 14Kxx, 14G35, 11F46, 11F80
@@
Keywords and phrases: modularity, abelian varieties, Siegel modular variety, Siegel modular forms, Galois
representation, zeta functions
@@ 1. The Modularity of an Elliptic Curve
We set $\Gamma_{1}:=SL(2,\mathbb{Z})$. For a positive integer $N$, we let $\Gamma(N),\ \Gamma_{1}(N)$ and $\Gamma_{0}(N)$ be the congruence subgroups of $\Gamma_{1}$ such that $\Gamma(N)\subset\Gamma_{1}(N)\subset\Gamma_{0}(N)\subset\Gamma_{1}.$ We refer to [6, pp. 13–14, p. 21] for the precise definitions and properties of $\Gamma(N),\ \Gamma_{1}(N)$ and $\Gamma_{0}(N)$. Let $\mathbb{H}_{1}$ be the Poincaré upper half plance. The quotient
@@
be the complex manifold which has a natural model $Y_{1}(N)/\mathbb{Q}$ (resp. $Y_{0}(N)/\mathbb{Q}$). We let $X_{1}(N)$ (resp. $X_{0}(N)$) be the smooth projective curve which contains $Y_{1}(N)$ (resp. $Y_{0}(N)$) as a dense Zariski open subset (cf. see [6, pp. 45–60]).
Let $S_{k}(N)$ be the space of cusp forms of weight $k\geq 1$ and level $N\geq 1$. Here $k$ and $N$ be positive integers. We recall that if $f\in S_{k}(N)$, it satisfies the following properties : (C1) $f((a\tau+b)(c\tau+d)^{-1})=(c\tau+d)^{k}f(\tau)$ for all $\begin{pmatrix}a&b\\ c&d\end{pmatrix}\in\Gamma_{1}(N)$ and $\tau\in\mathbb{H}_{1}$; (C2) $|f(\tau)|^{2}({\rm Im}\,\tau)^{k}$ is bounded in $\mathbb{H}_{1}$ and (C3) the Fourier expansion of $f(\tau)$ is given by
@@
For each prime $p\!\not|N$, we recall that the Hecke operator $T_{p}:S_{k}(N)\longrightarrow S_{k}(N)$ is defined by
@@
for any $\begin{pmatrix}a&b\\ c&d\end{pmatrix}\in\Gamma_{1}$ with $c\equiv 0\,({\rm mod}\,N)$ and $d\equiv p\,({\rm mod}\,N)$. We refer to [2, p. 844] or [6, pp. 170–171] for more detail. The Hecke operators $T_{p}\,(p\!\!\not|N)$ can be simultaneously diagonalized on $S_{k}(N)$ and a simultaneous eigenvector a Hecke eigenform or simply an eigenform.
Let $\lambda$ be a place of the algebraic closure $\bar{\mathbb{Q}}$ of $\mathbb{Q}$ in $\mathbb{C}$ lying over a rational integer $\ell$ and ¯${\bar{\mathbb{Q}}}_{\lambda}$ denote the algebraic closure of $\mathbb{Q}_{\ell}$ via $\lambda$. Let $G_{\mathbb{Q}}:={\rm Gal}({\bar{\mathbb{Q}}}/\mathbb{Q})$ be the absolute Galois group of $\mathbb{Q}$. It is well known that if $f\in S_{k}(N)$ is a normalized eigenform with $a_{1}(f)=1$, then there exists a unique continuous irreducible Galois representation
@@
such that $\rho_{f,\lambda}$ is unramified at $p$ for all primes $p\!\!\not|\,\ell N$ and
${\rm Tr}\left(\rho_{f,\lambda}({\rm Frob}_{p})\right)=a_{p}(f)\quad{\rm for\ any\ prime}\ p\!\!\not|\,\ell N.$
@@
$L(E,s):=\prod_{p|D}\left(\frac{1}{1-a_{p}(E)p^{-s}}\right)\prod_{p\not|D}\left(\frac{1}{1-a_{p}(E)\,p^{-s}+p^{1-2s}}\right).$
@@
be the Jacobian variety of the modular curve $X_{1}(N)$. Here $\Omega^{1}(X_{1}(N))$ denotes the complex vector space of holomorphic 1-forms on $X_{1}(N)$ and $W^{\vee}$ denotes the dual space of a complex vector space $W$. It is known that the following statements are equivalent : (a) $E$ is modular. (b) There is a non-constant holomorphic mapping $X_{1}(N)\longrightarrow E(\mathbb{C})$ for some positive
integer $N$. (c) There is a non-constant holomorphic mapping $J_{1}(N)\longrightarrow E(\mathbb{C})$ for some positive
integer $N$. (d) $\rho_{E,\ell}$ is modular for a prime $\ell$.
The above statements have been called the Taniyama-Shimura conjecture. We refer to [18] for the historical story of this conjecture. The implication (a) $\Longrightarrow$ (b) follows from a construction of Shimura [17] and a theorem of Faltings [7]. The implication (b) $\Longrightarrow$ (d) is due to Mazur [14]. The implication (d) $\Longrightarrow$ (a) follows from a theorem of Carayol [3]. The implication (c) $\Longrightarrow$ (b) is obvious. Wiles [21, 22] proved that a semistable eliptic curve over $\mathbb{Q}$ is modular by proving the statement (d). Thereafter Breuil, Conrad, Diamond and Taylor [2] proved that every eliptic curve over $\mathbb{Q}$ is modular.
Serre [16] conjectured the following : Serre’s Modularity Conjecture : Let $\overline{\rho}:G_{\mathbb{Q}}\longrightarrow GL(2,{\mathbb{F}})$ be a two-dimensional absolutely irreducible, continuous, odd representation of $G_{\mathbb{Q}}$. Here $\mathbb{F}$ is a finite field of characteristic $p$. Then $\overline{\rho}$ is modular, i.e., arises from (with respect to some fixed embedding $\imath:\overline{\mathbb{Q}}\hookrightarrow\overline{\mathbb{Q}_{p}})$ a newform $f$ of some weight $k\geq 2$ and level $N$ prime to $p$.
@@ 2. The Modularity of an Abelian Variety
$\alpha\cdot\Omega=(A\Omega+B)(C\Omega+D)^{-1},$
@@
$G/K\longrightarrow\mathbb{H}_{g},\qquad\alpha K\mapsto\alpha\!\cdot\!iI_{g},\quad\alpha\in G.$
@@
Let $\Gamma$ be a congruence subgroup of $\Gamma_{g}$ and let $k$ be a nonnegative integer $k$. A function $F:\mathbb{H}_{g}\longrightarrow\mathbb{C}$ is called a $\sf{Siegel\ modular\ form}$ of degree $g$ and weight $k$ with respect to $\Gamma$ if it satisfies the following conditions : (S1) $F(\Omega)$ is holomorphic on $\mathbb{H}_{g}$ ; (S2) $F(\gamma\cdot\Omega)=(C\Omega+D)^{k}F(\Omega)$   for all $\gamma=\begin{pmatrix}A&B\\ C&D\end{pmatrix}\in\Gamma$ and $\Omega\in\mathbb{H}_{g}$ ; (S3) $F(\Omega)$ is bounded in any domain $Y\geq Y_{0}>0$ in the case $g=1$.
We denote the space of all Sigel modular forms of degree $g$ and weight $k$ with respect to $\Gamma$ by $[\Gamma,k]$. We define the so-called $\sf{Siegel\ operator}$
@@
Then $\Phi_{g}$ is a well-defined linear mapping (cf, [13, pp. 187–189]). A Siegel modular form $F\in[\Gamma_{g},k]$ is called a $\sf{Siegel\ cusp\ form}$ if $\Phi_{g}(F)=0$ (cf. [13, p. 198]). We denote by $[\Gamma_{g},k]_{0}$ the space of all Siegel cusp forms in $[\Gamma_{g},k]$.
@@
where $T$ runs through all $g\times g$ half-integral semi-positive symmetric matrices. Here ${\rm Tr}(M)$ denotes the trace of a square matrix $M$. Following the Hecke’s method, Maass [13, pp. 202–210] associated with $F(\Omega)$ the Dirichlet series
@@
Let $F$ be a nonzero Siegel Hecke eigenform in $[\Gamma_{g},k]_{0}$. Let $\alpha_{p,0},\alpha_{p,1},\cdots,\alpha_{p,g}$ be the $p$-Satake parameters of $F$ at a prime $p$. We define the $\sf{local\ spinor\ zeta\ function}$ $Z_{F,p}(t)$ of $F$ at $p$ by
@@
Secondly one has the so-called $\sf{standard\ zeta\ function}$ $D_{F}(s)$ of a Siegel Hecke eigenform $F$ in $[\Gamma_{g},k]_{0}$ defined by
@@
(1) If $g=1$, the spinor zeta function $Z_{f}(s)$ of a Hecke eigenform $f$ in $S_{k}(\Gamma_{1})$ is nothing but the Hecke $L$-function $L(f,s)$ of $f$. (2) If $g=1$, the standard zeta function $D_{f}(s)$ of a Hecke eigenform $f(\tau)=\sum_{n=1}^{\infty}a(n)\,e^{2\pi in\tau}$ in $S_{k}(\Gamma_{1})$ has the following equation

### v2

@@ Abstract.
We introduce the concept of the modularity of an abelian variety defined over the rational number field extending the modularity of an elliptic curve. We discuss the modularity of an abelian variety over $\mathbb{Q}$. We conjecture that a simple abelian variety over $\mathbb{Q}$ is modular.
Subject Classification: Primary 14Kxx, 14G35, 11F46, 11F40, 11F80
@@
Keywords and phrases: modularity, abelian variety, Siegel modular variety, Galois representation,
Siegel Hecke eigenforms, zeta functions, Hecke algebra, $p$-Satake parameters.
1. Introduction
An elliptic curve $E$ over $\mathbb{Q}$ is said to be $\sf modular$} if the $L$-function $L(E,s)$ of $E$ equals the $L$-function $L(f,s)$ for some eigenform $f$, equivalently if $E$ has a finite covering by a modular curve of the form $X_{0}(N)$. At the Tokyo-Nikko conference held in 1955, Yutaka Taniyama (1927–1958) made a suggestion that every elliptic curve over $\mathbb{Q}$ is modular. At that time his suggestion was not clear and hence was not known in the mathematics community. In the early 1960’s, Goro Shimura (1930–2019) refined Taniyama’s suggestion through private conversations with a number of mathematicians and their efforts. In particular he discussed this subject with André Weil (1906–1998) seriously and intensively. Weil gave conceptual evidence for Taniyama’s suggestion in his famous paper [1] published in 1967. Through Weil’s paper, this suggestion was widely known as the so-called Shimura-Taniyama conjecture in the mathematics community.
The Shimura-Taniyama conjecture associates objects of representation theory to those of algebraic geometry. It states that the $L$-series of an elliptic curve over $\mathbb{Q}$ which measures the behaviour of the curve mod $p$ for all primes $p$, can be identified with an integral transform of the Fourier series defined from an eigenform.
In 1985 Frey [2] made the remarkable observation that the Shimura-Taniyama conjecture for a semistable elliptic curve over $\mathbb{Q}$ should imply Fermat’s Last Theorem. The precise mechanism relating the two was formulated by Serre [3] as the $\epsilon$-conjecture and this conjecture was proved by Ribet [4] in the summer of 1986. Ribet’s result only requires to prove the Shimura-Taniyama conjecture for a semistable elliptic curve over $\mathbb{Q}$ in order to reduce Fermat’s Last Theorem. As soon as Andrew Wiles learned Ribet’s result, he began to work the Shimura-Taniyama conjecture for a semistable elliptic curve over $\mathbb{Q}$ in the late summer of 1986. Here the semistabilty of an elliptic curve $E$ over $\mathbb{Q}$ is defined as follows :
Definition 1.1.
An elliptic curve $E$ over $\mathbb{Q}$ is said to be semistable at the prime $q$ if it is isomorphic to an elliptic curve $\tilde{E}$ over $\mathbb{Q}$ which $\tilde{E}$(mod $q$) is either nonsingular or has a node. An elliptic curve over $\mathbb{Q}$ is called semistable if it is semistable at every prime.
On June 21-23 in 1993, Wiles had given a series of lectures under the title, ”Elliptic curves, modular forms and Galois representations” at the Issac Newton Institute for Mathematical Sciences in Cambridge, England. In this last lecture on June 23, Wiles commented that he had proved a part of the Shimura-Taniyama conjecture to the effect that every semistable elliptic curve over $\mathbb{Q}$ is modular. With the aide of the works of Frey [2], Serre [3] and Ribet [4], his proof solves Fermat’s Last Theorem which had been unsolved for more than 350 years. The news spread out all over the world through the well-known newspapers and magazines because Fermat’s Last Theorem holds great fascination for amateurs and professionals alike. But in the fall that year it turned out that the proof of Wiles was incomplete and flawed. Precisely, his construction of the Euler system used to extend Flach’s method was not complete. He was very concerned with filling the gaps in his flawed proof. At that time he was completely isolated from outside. In January 1994, he proposed Dr. Richard Taylor, his former Ph.D. student in Cambridge University, UK to join him in the attempt to repair the Euler system argument. Wiles still was convinced that his method was correct. Dr. Taylor accepted his proposal and joined in that project. On September 19th in 1994, Wiles was quite convinced that his method was correct and his gap could be filled up. After he invited Dr. Taylor to Princeton again, he completed his proof with the aid of Dr. Taylor on October, 1994 and submitted his paper to Annals of Mathematics on October 14, 1994. Finally his paper was accepted and was published in May, 1995 (cf. [5, 6]). We refer to [7] for more interesting stories.
The aim of this article is to introduce the concept of the modularity of an abelian variety defined over the rational number field extending the modularity of an elliptic curve and to discuss the modularity of an abelian variety over $\mathbb{Q}$. An abelian variety is defined as follows in [8, p. 39] :
Definition 1.2.
An abelian variety $X$ is a complete [^1] This means, in particular, that it is irreducible. See the footnote in [8, p. 39]. algebraic variety over an algebraically closed field $K$ with a group law $m_{X}:X\times X\longrightarrow X$ such that both $m_{X}$ and the inverse map are morphisms of varieties. An abelian variety over $\mathbb{Q}$ is a connected and complex projective manifold that is also a group variety generalizing elliptic curves over $\mathbb{Q}$.
The paper is organized as follows. In Section 2, we briefly outline the modularity of an elliptic curve over $\mathbb{Q}$. Wiles proved the modularity of a semistable elliptic curve over $\mathbb{Q}$ leading to solve Fermat’s Last Theorem (cf. [5, 6]). In 2001, Breuil, Conrad, Diamond and Taylor proved that every elliptic curve over $\mathbb{Q}$ is modular (cf. [9]). In Section 3, we introduce the notion of the modularity of an abelian variety over $\mathbb{Q}$. We propose several conjectures and open problems. In Section 4 (Appendix), we describe and survey the Hecke algebra for a symplectic group and $p$-Satake parameters for the readers.
Notations: We denote by $\mathbb{Q},\,\mathbb{R}$ and $\mathbb{C}$ the field of rational numbers, the field of real numbers and the field of complex numbers respectively. We denote by $\mathbb{Z}$ and $\mathbb{Z}^{+}$ the ring of integers and the set of all positive integers respectively. For a prime $p$, we denote by $\mathbb{Q}_{p}$ the field of $p$-adic numbers and by $\mathbb{Z}_{p}$ the ring of $p$-adic integers. $\overline{\mathbb{Q}}$ (resp. $\overline{\mathbb{Q}_{p}}$) denotes the algebraic closure of $\mathbb{Q}$ (resp. $\mathbb{Q}_{p}$). $G_{\mathbb{Q}}={\rm Gal}(\overline{\mathbb{Q}}/\mathbb{Q})$ denotes the absolute Galois group of $\mathbb{Q}$. The symbol “:=” means that the expression on the right is the definition of that on the left. For two positive integers $k$ and $l$, $F^{(k,l)}$ denotes the set of all $k\times l$ matrices with entries in a commutative ring $F$. For a square matrix $M\in F^{(k,k)}$ of degree $k$, ${\rm Tr}(M)$} denotes the trace of $M$. For any $M\in F^{(k,l)},\ ^{t}\!M$ denotes the transpose of a matrix $M$. $I_{n}$ denotes the identity matrix of degree $n$. For $A\in F^{(k,l)}$ and $B\in F^{(k,k)}$, we set $B[A]=\,^{t}ABA$ (Siegel’s notation). For a number field $F$, we denote by ${\mathbb{A}}_{F}$ the ring of adeles of $F$. If $F=\mathbb{Q}$, the subscript will be omitted.
$Sp(2g,\mathbb{R})=\left\{\,M\in Sp(2g,\mathbb{R})\,|\ {}^{t}MJ_{g}M=J_{g}[M]=J_{g}\,\right\}$
denotes the symplectic group of degree $g$, where
$J_{g}=\begin{pmatrix}0&I_{g}\\ -I_{g}&0\end{pmatrix}$.
$\Gamma_{g}:=Sp(2g,\mathbb{Z})=\left\{\ \begin{pmatrix}A&B\\ C&D\end{pmatrix}\in Sp(2g,\mathbb{R})\Big|\ A,B,C,D\ \rm{integral}\,\right\}$
denotes the Siegel modular group of degree $g$. For a prime $\ell$ and a positive integer $n$, ${\mathbb{F}}_{\ell^{n}}$ denotes the field with $\ell^{n}$ elements. The algebraic closure $\overline{\mathbb{F}}_{\ell}$ of $\mathbb{F}$$_{\ell}$ is given by
$\overline{\mathbb{F}}_{\ell}=\bigcup_{n\geq 1}{\mathbb{F}}_{\ell^{n}}.$
If $p$ is prime and $\lambda$ is a prime ideal dividing $p$ in the ring of integers in $\overline{\mathbb{Q}}$, there exist a filtration
$I_{\lambda}\subset D_{\lambda}\subset G_{\mathbb{Q}},$
where the decomposition group $D_{\lambda}$ and the inertia group $I_{\lambda}$ are defined respectively by
$D_{\lambda}:=\{\,\sigma\in G_{\mathbb{Q}}\,|\,\sigma(\lambda)=\lambda\,\}$
and
$I_{\lambda}:=\{\,\sigma\in D_{\lambda}\,|\,\sigma(x)\equiv x\,({\rm mod}\,\lambda)\ {\rm{for\ all\ algebraic\ integers}}\ x\,\}.$
Then there are natural identifications
$D_{\lambda}\cong{\rm Gal}(\overline{\mathbb{Q}}_{p}/\mathbb{Q}_{p})\quad{\rm and}\quad D_{\lambda}/I_{\lambda}\cong{\rm Gal}(\overline{\mathbb{F}}_{p}/\mathbb{F}_{p}).$
We denote by ${\rm Frob}_{\lambda}\in D_{\lambda}/I_{\lambda}$ the inverse image of the canonical generator $x\mapsto x^{p}$ of ${\rm Gal}(\overline{\mathbb{F}}_{p}/{\mathbb{F}}_{p}).$ If $\lambda^{\prime}$ is another prime lying above $p$, then $\lambda^{\prime}=\sigma(\lambda)$ for some $\sigma\in G_{\mathbb{Q}}$ and
$D_{\lambda^{\prime}}=\sigma D_{\lambda}\sigma^{-1},\quad I_{\lambda^{\prime}}=\sigma I_{\lambda}\sigma^{-1}\quad{\rm and}\quad{\rm Frob}_{\lambda^{\prime}}=\sigma\,{\rm Frob}_{\lambda}\,\sigma^{-1}.$
Since we will care about these objects only up to conjugation, we will write $D_{\lambda}$ and $I_{\lambda}$. Now we will write ${\rm Frob}_{p}$ for any representative of a ${\rm Frob}_{\lambda}$. If $\rho$ is a representation of $G_{\mathbb{Q}}$ which is unramified at $p$, then ${\rm Tr}(\rho({\rm Frob}_{p}))$ and $\det(\rho({\rm Frob}_{p}))$ are well defined, that is, are independent of the choice of $\lambda$.
@@ 2. The Modularity of an Elliptic Curve
We set $\Gamma_{1}:=SL(2,\mathbb{Z})$. For a positive integer $N$, we let $\Gamma(N),\ \Gamma_{1}(N)$ and $\Gamma_{0}(N)$ be the congruence subgroups of $\Gamma_{1}$ such that $\Gamma(N)\subset\Gamma_{1}(N)\subset\Gamma_{0}(N)\subset\Gamma_{1}.$ We refer to [10, pp. 13–14, p. 21] for the precise definitions and properties of $\Gamma(N),\ \Gamma_{1}(N)$ and $\Gamma_{0}(N)$. Let $\mathbb{H}_{1}$ be the Poincaré upper half plance. The quotient
@@
be the complex manifold which has a natural model $Y_{1}(N)/\mathbb{Q}$ (resp. $Y_{0}(N)/\mathbb{Q}$). We let $X_{1}(N)$ (resp. $X_{0}(N)$) be the smooth projective curve which contains $Y_{1}(N)$ (resp. $Y_{0}(N)$) as a dense Zariski open subset (cf. see [10, pp. 45–60]).
Let $S_{k}(N)$ be the space of cusp forms of weight $k\geq 1$ and level $N\geq 1$. Here $k$ and $N$ be positive integers. We recall that if $f\in S_{k}(N)$, it satisfies the following properties :
(C1) $f((a\tau+b)(c\tau+d)^{-1})=(c\tau+d)^{k}f(\tau)$ for all $\begin{pmatrix}a&b\\ c&d\end{pmatrix}\in\Gamma_{1}(N)$ and $\tau\in\mathbb{H}_{1}$;
(C2) $|f(\tau)|^{2}({\rm Im}\,\tau)^{k}$ is bounded in $\mathbb{H}_{1}$ and
(C3) the Fourier expansion of $f(\tau)$ is given by
@@
[…]
@@
[…]
Let $\lambda$ be a place of the algebraic closure $\bar{\mathbb{Q}}$ of $\mathbb{Q}$ in $\mathbb{C}$ lying over a rational integer $\ell$ and ${\bar{\mathbb{Q}}}_{\lambda}$ denote the algebraic closure of $\mathbb{Q}_{\ell}$ via $\lambda$. Let $G_{\mathbb{Q}}:={\rm Gal}({\bar{\mathbb{Q}}}/\mathbb{Q})$ be the absolute Galois group of $\mathbb{Q}$. It is well known that if $f\in S_{k}(N)$ is a normalized eigenform with $a_{1}(f)=1$, then there exists a unique continuous irreducible Galois representation
@@
[…]
@@
[…]
@@
be the Jacobian variety of the modular curve $X_{1}(N)$. Here $\Omega^{1}(X_{1}(N))$ denotes the complex vector space of holomorphic 1-forms on $X_{1}(N)$ and $W^{\vee}$ denotes the dual space of a complex vector space $W$. It is known that the following statements are equivalent :
(a) $E$ is modular.
(b) There is a non-constant holomorphic mapping $X_{1}(N)\longrightarrow E(\mathbb{C})$ for some pos-
itive integer $N$.
(c) There is a non-constant holomorphic mapping $J_{1}(N)\longrightarrow E(\mathbb{C})$ for some pos-
itive integer $N$.
(d) $\rho_{E,\ell}$ is modular for a prime $\ell$.
The above statements have been called the Taniyama-Shimura conjecture. We refer to [15] for the historical story of this conjecture. The implication (a) $\Longrightarrow$ (b) follows from a construction of Shimura [11] and a theorem of Faltings [16]. The implication (b) $\Longrightarrow$ (d) is due to Mazur [17]. The implication (d) $\Longrightarrow$ (a) follows from a theorem of Carayol [18]. The implication (c) $\Longrightarrow$ (b) is obvious. Wiles [5, 6] proved that a semistable elliptic curve over $\mathbb{Q}$ is modular by proving the statement (d). Thereafter Breuil, Conrad, Diamond and Taylor [9] proved that every elliptic curve over $\mathbb{Q}$ is modular.
Serre [3] conjectured the following :
Serre’s Modularity Conjecture : Let $\overline{\rho}:G_{\mathbb{Q}}\longrightarrow GL(2,{\mathbb{F}})$ be a two-dimensional absolutely irreducible, continuous, odd representation of $G_{\mathbb{Q}}$. Here $\mathbb{F}$ is a finite field of characteristic $p$. Then $\overline{\rho}$ is modular, i.e., arises from (with respect to some fixed embedding $\imath:\overline{\mathbb{Q}}\hookrightarrow\overline{\mathbb{Q}_{p}})$ a newform $f$ of some weight $k\geq 2$ and level $N$ prime to $p$.
@@ 3. The Modularity of an Abelian Variety
(3.1) $\alpha\cdot\Omega=(A\Omega+B)(C\Omega+D)^{-1},$
@@
$G/K\longrightarrow\mathbb{H}_{g},\qquad\alpha K\longmapsto\alpha\!\cdot\!iI_{g},\quad\alpha\in G.$
@@
Let $\Gamma$ be a congruence subgroup of $\Gamma_{g}$ and let $k$ be a nonnegative integer $k$. A function $F:\mathbb{H}_{g}\longrightarrow\mathbb{C}$ is called a $\sf{Siegel\ modular\ form}$ of degree $g$ and weight $k$ with respect to $\Gamma$ if it satisfies the following conditions :
(S1) $F(\Omega)$ is holomorphic on $\mathbb{H}_{g}$ ;
(S2) $F(\gamma\cdot\Omega)=(C\Omega+D)^{k}F(\Omega)$ for all $\gamma=\begin{pmatrix}A&B\\ C&D\end{pmatrix}\in\Gamma$ and $\Omega\in\mathbb{H}_{g}$ ;
(S3) $F(\Omega)$ is bounded in any domain $Y\geq Y_{0}>0$ in the case $g=1$.
We denote the space of all Siegel modular forms of degree $g$ and weight $k$ with respect to $\Gamma$ by $\Gamma,k]$.
We define the so-called $\sf{Siegel\ operator}$
@@
Then $\Phi_{g}$ is a well-defined linear mapping (cf, [21, pp. 187–189]). A Siegel modular form $F\in[\Gamma_{g},k]$ is called a $\sf{Siegel\ cusp\ form}$ if $\Phi_{g}(F)=0$ (cf. [21, p. 198]). We denote by $\Gamma_{g},k]_{0}$ the space of all Siegel cusp forms in $\Gamma_{g},k]$.
@@
where $T$ runs through all $g\times g$ half-integral semi-positive symmetric matrices. Here ${\rm Tr}(M)$ denotes the trace of a square matrix $M$. Following Hecke’s method, Maass [21, pp. 202–210] associated with $F(\Omega)$ the Dirichlet series
@@
Let $F$ be a nonzero Siegel Hecke eigenform in $\Gamma_{g},k]_{0}$. Let $\alpha_{p,0},\alpha_{p,1},\cdots,$
$\alpha_{p,g}$ be the $p$-Satake parameters of $F$ at a prime $p$ (cf. see Section 4). We define the $\sf{local\ spinor\ zeta\ function}$ $Z_{F,p}(t)$ of $F$ at $p$ by
@@
Secondly one has the so-called $\sf{standard\ zeta\ function}$ $D_{F}(s)$ of a Siegel Hecke eigenform $F$ in $\Gamma_{g},k]_{0}$ defined by
@@
(1) If $g=1$, the spinor zeta function $Z_{f}(s)$ of a Hecke eigenform $f$ in $S_{k}(\Gamma_{1})$ is nothing but the Hecke $L$-function $L(f,s)$ of $f$.
(2) If $g=1$, the standard zeta function $D_{f}(s)$ of a Hecke eigenform $f(\tau)=\sum_{n=1}^{\infty}a(n)\,e^{2\pi in\tau}$ in $S_{k}(\Gamma_{1})$ has the following equation